OpenStudy (anonymous):

The coefficient of x in the expansion of (x+1/ax^2)^7 is 7/3. Find the possible values of a.

OpenStudy (anonymous):

are there options?

OpenStudy (anonymous):

No,unfortunately not

OpenStudy (anonymous):

ok, after expansion the x coefficient is (21/a^2) \[\frac{7}{3} = \frac{21}{a^{2}}\] \[a^{2} = \frac{21*3}{7} = 9\] \[a = \pm 3\] the term that has x to the power of 1 should be 21/a^2 you need to simplify the terms first if you simplify that term, you get --> 7/(a^6*x^11) Use term 21x^5(1/(ax^2))^2 = (21x^5)/(a^2*x^4) = 21x/a^2